# Weak lensing of bright standard sirens: prospects for $\sigma_{8}$

### v1

@@ III Hubble diagram reconstruction
where $V_{c}$ denotes the comoving volume. The power-law index $2.7$ is motivated by studies of the star formation rate density [47] and is consistent with the LVK results [48]. The redshifts of the BMBHs are drawn from a population generated using a merger rate derived from the extended Press-Schechter formalism combined with the MBH-galaxy co-evolution model of [49]. We restrict the BNS population to $z_{s}<2$ and the BMBH population to $z_{s}<10$ (see e.g. [43]). For both populations, we assume negligible uncertainty in the redshift measurements. For the apparent luminosity distances, we adopt a relative measurement uncertainty of $\sigma_{D_{L}}=0.03D_{L}$ for the BNSs and $\sigma_{D_{L}}=0.003D_{L}$ for the BMBHs. The catalogues are generated by sampling source redshifts from the respective redshift distributions and drawing a lensing magnification factor $\mu$ from the corresponding magnification PDFs. The resulting benchmark catalogues are shown in Fig. 4.
@@
In Fig. 5 we show the posterior distributions obtained for ET and LISA using the benchmark catalogues from Fig. 4. The posteriors are consistent with the fiducial values (dashed lines), demonstrating that the inference is unbiased. ET provides significantly tighter constraints than LISA, despite LISA observing more distant sources. This difference is driven by the larger expected catalogue size for ET. In the $\Omega_{M}-h$ panel, a pronounced negative correlation is evident, indicating that lower Hubble constant values can be partially compensated by higher matter density. In contrast, $\sigma_{8}$ shows much weaker correlations with the other parameters. Quantitatively, we find that ET can measure $\sigma_{8}$ with an accuracy of 10% and LISA with an accuracy of $30\%$, while their combination would yield an accuracy of $8\%$. The prospects for $h$ and $\Omega_{M}$ are similar to previous studies (e.g. [16]).
IV Conclusions
As GW detector sensitivities and electromagnetic follow-up capabilities continue to improve, bright standard sirens will become an increasingly important probe of cosmology. We have shown that weak-lensing-induced scatter in the luminosity distance-redshift relation encodes valuable information about the distribution of cosmic structures, enabling constraints on parameters that do not directly affect the mean Hubble diagram. Focusing on the amplitude of matter fluctuations, $\sigma_{8}$, we find that ET could achieve $10\%$ measurement with a sample of 300 neutron star binaries with identified electromagnetic counterparts. At lower GW frequencies, observations by LISA of massive black hole binaries provide a complementary avenue, yielding constraints at the $30\%$ level with a smaller sample of only 12 events.
Our results highlight that weak lensing transforms bright sirens from purely geometric probes into tracers of cosmic structures, opening a new channel for studying cosmological perturbations with GWs. While the accuracy of constraints inferred from CMB observations [41] will probably not be reached, weak lensing of bright sirens provides a more direct probe of low- to intermediate-redshift matter fluctuations. Furthermore, this approach is complementary to traditional probes such as galaxy clustering and weak-lensing surveys [50].
@@ IV Conclusions
The analysis can be readily extended to additional cosmological parameters beyond those considered in this study. In particular, including parameters such as the spectral tilt and spatial curvature would lead to more robust and realistic forecasts [20]. Furthermore, exploring scenarios beyond the standard cold dark matter paradigm would be of interest. Since alternative dark matter models generally predict different abundances and internal properties of cosmic structures, the resulting weak-lensing-induced scatter of bright sirens could provide novel and complementary constraints on dark matter properties. For example, warm or fuzzy dark matter suppresses small-scale structures [51, 52], thereby reducing the weak-lensing-induced scatter, whereas compact dark matter objects would enhance the high-magnification tail [53].
Acknowledgments – We are grateful to G. Congedo, K. Kainulainen and J. Urrutia for valuable discussions. This work was supported by the European Union’s Horizon Europe program under the Marie Skłodowska-Curie grant agreement No. 101065736, the Estonian Research Council grants TARISTU24-TK3 and TARISTU24-TK10, and the Center of Excellence program TK202 of the Estonian Ministry of Education and Research.

### v2

@@ 3 Hubble diagram reconstruction
where $V_{c}$ denotes the comoving volume. The power-law index $2.7$ is motivated by studies of the star formation rate density (Madau & Dickinson, 2014) and is consistent with the LVK results (Abac et al., 2025b). The redshifts of the BMBHs are drawn from a population generated using a merger rate derived from the extended Press-Schechter formalism combined with the MBH-galaxy co-evolution model of Urrutia et al. (2025). We restrict the BNS population to $z_{s}<2$ and the BMBH population to $z_{s}<10$ (see e.g. Mangiagli et al. (2025)). For both populations, we assume negligible uncertainty in the redshift measurements. For the apparent luminosity distances, we adopt a relative measurement uncertainty of $\sigma_{D_{L}}=0.03D_{L}$ for the BNSs and $\sigma_{D_{L}}=0.003D_{L}$ for the BMBHs.666 These measurement uncertainties are intentionally optimistic compared to population averaged forecasts, presented for BNSs, for example, by Iacovelli et al. (2022); Loffredo et al. (2025), and for BMBHs by Mangiagli et al. (2025), as they correspond to events for which an EM counterpart is identified. In such cases, information on the sky location and binary inclination from the EM measurement can substantially reduce the distance-inclination degeneracy in the GW signal (Holz & Hughes, 2005; Shah et al., 2013; Hotokezaka et al., 2019; Chen et al., 2019; Nakar & Piran, 2021; de Souza & Sturani, 2023). The catalogues are generated by sampling source redshifts from the respective redshift distributions and drawing a lensing magnification factor $\mu$ from the corresponding magnification PDFs. The resulting benchmark catalogues are shown in Fig. 4.
@@
In Fig. 5 we show the posterior distributions obtained for ET and LISA using the benchmark catalogues from Fig. 4. The posteriors are consistent with the fiducial values (dashed lines), demonstrating that the inference is unbiased. ET provides significantly tighter constraints than LISA, despite LISA observing more distant sources. This difference is driven by the larger expected catalogue size for ET. In the $\Omega_{M}-h$ panel, a pronounced negative correlation is evident, indicating that lower Hubble constant values can be partially compensated by higher matter density. In contrast, $\sigma_{8}$ shows much weaker correlations with the other parameters. Quantitatively, we find that ET can measure $\sigma_{8}$ with an accuracy of 10% and LISA with an accuracy of $30\%$, while their combination would yield an accuracy of $8\%$. The prospects for $h$ and $\Omega_{M}$ are similar to previous studies (e.g. Mpetha et al. (2024)).
4 Conclusions
As GW detector sensitivities and EM follow-up capabilities continue to improve, bright standard sirens will become an increasingly important probe of cosmology. We have shown that weak-lensing-induced scatter in the luminosity distance-redshift relation encodes valuable information about the distribution of cosmic structures, enabling constraints on parameters that do not directly affect the mean Hubble diagram. Focusing on the amplitude of matter fluctuations, $\sigma_{8}$, we find that ET could achieve $10\%$ measurement with a sample of 300 neutron star binaries with identified EM counterparts. At lower GW frequencies, observations by LISA of massive black hole binaries provide a complementary avenue, yielding constraints at the $30\%$ level with a smaller sample of only 12 events.
Our results highlight that weak lensing transforms bright sirens from purely geometric probes into tracers of cosmic structures, opening a new channel for studying cosmological perturbations with GWs. While the precision of constraints from CMB observations (Aghanim et al., 2020) or from galaxy clustering and weak-lensing surveys (Abbott et al., 2025) is unlikely to be reached, weak lensing of bright sirens provides a complementary probe of $\sigma_{8}$. Moreover, standard candles and standard sirens are uniquely sensitive to small-scale matter fluctuations that are largely inaccessible to conventional CMB or galaxy weak-lensing observations. In this work, we have focused on $\sigma_{8}$, but we plan to use bright-siren lensing to directly constrain the small-scale matter power spectrum.
@@ 4 Conclusions
The analysis can be readily extended to additional cosmological parameters beyond those considered in this study. In particular, including parameters such as the spectral tilt and spatial curvature would lead to more robust and realistic forecasts (Congedo & Taylor, 2019). Furthermore, exploring scenarios beyond the standard cold dark matter paradigm would be of interest. Since alternative dark matter models generally predict different abundances and internal properties of cosmic structures, the resulting weak-lensing-induced scatter of bright sirens could provide novel and complementary constraints on dark matter properties. For example, warm or fuzzy dark matter suppresses small-scale structures (Hu et al., 2000; Bode et al., 2001), thereby reducing the weak-lensing-induced scatter, whereas compact dark matter objects would enhance the high-magnification tail (Zumalacarregui & Seljak, 2018).
Acknowledgments
We are grateful to G. Congedo, K. Kainulainen and J. Urrutia for valuable discussions. This work was supported by the European Union’s Horizon Europe program under the Marie Skłodowska-Curie grant agreement No. 101065736, the Estonian Research Council grants TARISTU24-TK3 and TARISTU24-TK10, and the Center of Excellence program TK202 of the Estonian Ministry of Education and Research.
Data availability
The data generated for this article are available upon request from the author. The code used to generate the data is publicly available at https://github.com/vianvask/halos.